### v1

# OE_^ OmOE_^ OmOE_^ Omm!OE_^ mmmmOE_^

Abstract
Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities across diverse tasks, yet they lag significantly behind humans in spatial reasoning. We investigate this gap through Transformation-Driven Visual Reasoning (TVR), a challenging task requiring identification of object transformations across images under varying viewpoints. While traditional Supervised Fine-Tuning (SFT) fails to generate coherent reasoning paths in cross-view settings, sparse-reward Reinforcement Learning (RL) suffers from inefficient exploration and slow convergence. To address these limitations, we propose STAR-R1, a novel framework that integrates a single-stage RL paradigm with a fine-grained reward mechanism tailored for TVR. Specifically, STAR-R1 rewards partial correctness while penalizing excessive enumeration and passive inaction, enabling efficient exploration and precise reasoning. Comprehensive evaluations demonstrate that STAR-R1 achieves state-of-the-art performance across all 11 metrics, outperforming SFT by 23% in cross-view scenarios. Further analysis reveals STAR-R1’s anthropomorphic behavior and highlights its unique ability to compare all objects for improving spatial reasoning. Our work provides critical insights in advancing the research of MLLMs and reasoning models. The codes, model weights, and data will be publicly available at https://github.com/zongzhao23/STAR-R1.
1 Introduction
Multimodal Large Language Models (MLLMs) bai2023qwenvl ; wang2024qwen2vl ; bai2025qwen25vl ; chen2024internvl ; chen2024internvl1.5 ; team2024internvl2 ; chen2024expanding2.5 ; yao2024minicpm ; agrawal2024pixtral ; Emu2 ; wang2024emu3 ; chen2025janus have demonstrated remarkable progress in recent years, excelling at tasks such as visual question answering fu2024mme ; yu2023mm ; yue2024mmmu , text-to-image generation ghosh2023geneval ; hu2024ella , and video understanding li2024mvbench ; zhou2024mlvu ; wang2024lvbench ; fu2024video ; chen2024we . Despite these advances, existing MLLMs still struggle with spatial reasoning mayer2025ivispar ; tang2025lego —a fundamental aspect of human intelligence. Given this critical gap in emulating human-like spatial cognition, it becomes imperative to develop a targeted framework for assessing such aspect. To this end, this paper takes the challenging task of Transformation Driven Visual Reasoning (TVR) hong2021transformation as a starting point to investigate MLLMs’ capabilities in spatial understanding. By comparing two given images containing multiple objects, TVR involves a series of continuous thinking and reasoning steps to derive the correct answer, including identifying which object is transformed and what transformation is applied to this object. Moreover, objects in the two images may be presented in different viewpoints, which further complicates the task and poses difficulties for existing MLLMs. Our experiments show that, the powerful closed-source commercial model GPT-4o hurst2024gpt achieves only 23.5% accuracy on this task, further demonstrating that this problem poses a significant challenge to existing MLLMs.
Figure 1: Comparison of addressing the TVR task by STAR-SFT and STAR-R1. STAR-SFT finetunes MLLMs with supervised instructions, while STAR-R1 employs RL-guided thinking.
While Supervised Fine-Tuning (SFT) can improve MLLMs’ performance on TVR, this approach remains insufficient. Since successfully solving TVR requires step-by-step Chain-of-Thought (CoT) wei2022chain reasoning, merely replicating annotations of final object and attribute changes fails to yield robust performance, particularly in view-shifting scenarios. Recently, training LLM by Reinforcement Learning (RL), exemplified by DeepSeek-R1-Zero guo2025deepseek , has demonstrated significant advantages in generating the correct CoT, offering a promising solution for the TVR task. However, the naive GRPO method shao2024deepseekmath in DeepSeek-R1-Zero, designed for tasks like Mathematical Problem Solving (MPS) cobbe2021training ; hendrycks2021measuring ; hendrycks2020measuring ; zhong2023agieval ; wei2023cmath , uses sparse reward based on the binary verification signal. When directly adapted to the TVR task, it results in low exploration efficiency and slow convergence. In the MPS task, the final answer, such as the option of the multiple-choice question or the value of the fill-in-the-blank question, has only two possibilities: correct or incorrect, without intermediate state of partial correctness. In contrast, the final answer in the TVR task is a sequence composed of multiple objects with multiple transformations. During early training, the model often produces partially correct responses; for instance, it may correctly identify the object whose attribute has changed but fail to specify the exact transformation applied. Under a binary reward setting, the model rarely receives positive feedback, hindering its ability to determine effective exploration directions.
To address these challenges, we propose to integrate dense reward plus punishment into GRPO, thereby constructing a new method dubbed Spacial TrAnsformation Reasoning with R1 (STAR-R1). STAR-R1 is built upon two core designs. On the one hand, it introduces dense rewards by accounting for partial answer correctness, progressively assigning fine-grained rewards to three-level answers. Specifically, for the first-level answer, the model identifies the objects with modified attributes, receiving a minimal reward. For the second-level answer, it correctly predicts the changed attributes, earning a medium reward. For the third-level answer, the model accurately identifies the full object-attribute-transformation triplet, receiving the highest reward. On the other hand, STAR-R1 penalizes the model for either excessively enumerating object-attribute-change triplets to maximize rewards or remaining inactive to avoid incorrect predictions, ensuring active exploration without shortcuts. This mechanism design instills into the model the learning philosophy that prioritizes attempting potentially incorrect answers over overlooking possible correct ones.
Through rigorous data filtering and sampling pipeline, we select training/test samples from the original TRANCE dataset (hong2021transformation, ). To comprehensively evaluate existing MLLMs on the TVR task, our test set includes samples whose initial and final images are captured either from the same view (In-Domain) or from more challenging different views (Out-of-Domain), totaling 4.5K samples. Extensive evaluations on both closed-source commercial and open-source MLLMs demonstrate that the proposed STAR-R1 not only significantly outperforms other models across all types of attribute transformations, but also exceeds the performance of SFT methods under the same training settings. Notably, in Out-of-Domain (OOD) tasks, STAR-R1 achieves a substantial 23% performance advantage over SFT-based approaches. Furthermore, through in-depth analysis of experimental results, we reveal STAR-R1’s anthropomorphic behavior and identify the core reason for RL’s superiority: it majorly compares all objects between the initial and final scenes, thereby improving reasoning in OOD scenarios. We also observe an intriguing phenomenon where the model’s response length initially decreases and then increases during training, indicating the specificity of the TVR task compared to other common reasoning tasks. Our contributions are summarized as follows:
  • We propose STAR-R1, a multimodal reasoning model trained via a single-stage pure RL paradigm akin to DeepSeek-R1-Zero. By integrating a customized reward function for RL optimization, STAR-R1 exhibits superior reasoning capabilities in spatial understanding.
  • In the TVR task, STAR-R1 achieves the best performance across all eleven evaluation metrics. Moreover, in the cross-view OOD task setting, it delivers a 23% TAcc gain over SFT, validating the R1-Zero paradigm’s potential in unlocking reasoning capabilities of the TVR task.
  • Through in-depth analysis of experimental results, we uncover STAR-R1’s anthropomorphic behavior and elucidate the mechanistic advantages of RL over SFT. We also systematically investigate the dynamic changes in response length. These findings provide unique insights for advancing MLLMs in spatial reasoning.
2 Related Work
Multimodal Large Language Model. MLLMs have made rapid advancements recently, achieving significant improvements in various image fu2024mme ; mathew2021docvqa and video understanding li2024mvbench ; mangalam2023egoschema tasks. The early LLaVA series models liu2023llava ; liu2024llava-1.5 ; li2024llava-next ; li2024llava-onevision are typical representatives of MLLMs, which use a simple MLP as the projectioner to map the CLIP-encoded radford2021learning ; sun2023eva-clip visual features into the semantic space of Large Language Models (LLMs) touvron2023llama ; touvron2023llama2 ; grattafiori2024llama3 ; bai2023qwen ; yang2024qwen2.5 , enabling the visual-language joint understanding. This architecture is subsequently widely adopted by more powerful models such as the QwenVL series bai2023qwenvl ; wang2024qwen2vl ; bai2025qwen25vl and InternVL series chen2024internvl ; chen2024internvl1.5 ; team2024internvl2 , significantly narrowing the performance gap between open-source MLLMs and commercial-grade closed-source benchmark models like GPT-4o hurst2024gpt and Gemini series team2024gemini . Despite these advancements, current MLLMs still struggle with complex reasoning tasks, especially those involving spatial reasoning mayer2025ivispar ; tang2025lego . To this end, we treat the TVR task as a pivotal point, concentrating on exploring effective methods for enhancing the spatial reasoning capabilities of MLLMs.
Multimodal Reasoning Model. Research on the reasoning capabilities of MLLMs has attracted significant attention in recent years. Early studies primarily employed Chain-of-Thought (CoT) fine-tuning to enhance model reasoning abilities (xu2024llava, ; xu2025redstar, ), while the recent success of DeepSeek-R1 (guo2025deepseek, ) has promoted the application of rule-based Reinforcement Learning in this field (chen2025r1v, ; shen2025vlm, ). Notably, MM-Eureka (meng2025mm, ) adopts an online filtering strategy and a two-stage multimodal RL training framework to achieve significant improvements in mathematical reasoning tasks; LMM-R1 (peng2025lmm, ), which enhances multimodal reasoning capabilities through a two-phase training paradigm combining text-only RL and multimodal RL; and Video-R1 (feng2025video, ), which achieves breakthroughs in video reasoning tasks via a two-stage approach involving supervised fine-tuning (SFT) followed by RL training. However, existing methods exhibit several limitations: (1) most rely on naive GRPO strategies, (2) they predominantly adopt two-stage training paradigms, and (3) there is a lack of task-specific customization and in-depth exploration of pure RL training (e.g., like DeepSeek-R1-Zero). To address these gaps, we introduces a tailored rule-based RL strategy for the TVR task and proposes a single-stage, pure RL training paradigm. Our approach not only significantly enhances multimodal reasoning performance but also provides novel theoretical insights into RL training mechanisms.
Figure 2: The STAR-R1 Framework. Given initial/final images and a question, STAR-R1 establishes object correspondences by systematically comparing all objects across both images. It generates Step-by-step reasoning in <think></think> tags before delivering the final answer in <answer></answer> tags. We demonstrate the reward function of STAR-R1 using a partially correct case study. Correct/incorrect answers are shown in blue and red respectively.
3 Our Method: STAR-R1
3.1 Task Formulation
As illustrated in Fig. 2, the TVR task is formally defined in this way: given an initial image $I^{\text{initial}}$, a final image $I^{\text{final}}$, and a question $Q$, the model $\phi$ is required to generate a sequence of transformations $T$ as the output answer. Formally, this generation process can be expressed as:
$T=\phi(I^{\text{initial}},I^{\text{final}},Q).$ (1)
Here, $T=\{t_{1},t_{2},\dots,t_{n}\}$ denotes a sequence of transformations. To avoid the enormous description space introduced by pixel-level transformations, we adopt the methodology from the previous work (hong2021transformation, ), defining a single-attribute transformation of a single object as an atomic operation. In form, $t_{i}=(\text{index}_{i},\text{attribute}_{i},\text{value}_{i})$, which indicates modifying $\text{attribute}_{i}$ to $\text{value}_{i}$, of the object with $\text{index}_{i}$. Given the multi-step nature of transformations, MLLMs must meticulously analyze fine-grained differences between the initial and final images and perform step-by-step reasoning to derive the correct transformation sequence.
3.2 The Design of Reward Function
Reward function, serves as direct feedback from the environment and plays a pivotal role in reinforcement learning. A sparse reward space can lead to inefficient exploration by the model, often requiring extended training periods to discover optimal decision paths. An excessively dense reward, on the contrary, may induce the model to adopt shortcut strategies for reward hacking. To facilitate active and efficient exploration of the solution space, our STAR-R1 integrates dense reward and punishment mechanisms into the original GRPO framework (shao2024deepseekmath, ). Particularly, we incorporate not only format rewards for structural constraints but also dense reward and punishment mechanisms to provide precise reward signals. The design principles of these two reward components are as follows.
Format Reward The format reward enforces structural constraints on model outputs, mandating that reasoning processes should appear within <think></think> tags while final answers must be enclosed in <answer></answer> tags. Compliance of this formate yields a reward $R^{\text{format}}$ of 1, whereas violations result in 0.
Accuracy Reward In existing studies, Reinforcement Learning with Verifiable Rewards (RLVR) primarily employs binary verification signals (correct or incorrect), which can be easily computed by simple rule-based verifiers. This paradigm has been validated in multiple studies and has demonstrated promising results across various scenarios  (xie2025logic, ; meng2025mm, ). However, in the TVR task, directly applying binary verification signals—the model only receives a reward if all transformations are correct—would lead to inefficient exploration and suboptimal model performance. Therefore, our model STAR-R1 first refines the reward granularity to each individual transformation and then decomposes the reward into each component of the transformation. Additionally, it carries out punishment mechanisms to guide the model toward correct and proactive exploration, ensuring both efficient exploration and high prediction accuracy.
For any transformation $t_{i}=(\text{index}_{i},\text{attribute}_{i},\text{value}_{i})$ in the model’s output set $T=\{t_{1},t_{2},\dots,t_{n}\}$, if its object index, attribute, and value exactly match a corresponding $\hat{t}_{i}$ in the ground-truth transformations $\hat{T}=\{\hat{t}_{1},\hat{t}_{2},\dots,\hat{t}_{n}\}$, the model is deemed to have perfectly inferred the transformation and thus receives the highest reward ($\text{reward}=5.0$). If only the object index and attribute match but the value is incorrect, the model demonstrates accurate identification and association capabilities, which are critical for forming correct reasoning chains; hence, a partial reward ($\text{reward}=1.5$) is granted. A match solely in the object index reflects preliminary identification ability, and a minimal reward ($\text{reward}=0.5$) is provided to encourage iterative model improvement. The positive reward described above can be formally defined as:
$R^{\text{pos}}(t_{i})=\begin{cases}5.0,&\text{if all $\text{index}_{i}$, $\text{attribute}_{i}$, and $\text{value}_{i}$ match;}\\ 1.5,&\text{if $\text{index}_{i}$ and $\text{attribute}_{i}$ match;}\\ 0.5,&\text{if only $\text{index}_{i}$ matches.}\end{cases}$ (2)
By accumulating the rewards of all $t_{i}$s, we obtain $R^{\text{pos}}=\sum^{n}_{i=1}R^{\text{pos}}(t_{i})$.
Building upon the positive reward design, we further introduce a dual punishment mechanism to effectively constrain the model’s exploratory behavior. Relying solely on positive rewards may lead the model to engage in reward hacking by exhaustively enumerating all possible $(\text{index}_{i},\text{attribute}_{i},\text{value}_{i})$ triplets. To mitigate this, after positive reward calculation, we further examine the predictions to verify whether the attribute transformations are inconsistent with the final image state. Each mistaken prediction incurs a punishment ($\text{reward}=-1.0$), yielding a total reward as $-n_{\text{miss}}$ for $n_{\text{miss}}$ mistaken predictions. Furthermore, to encourage comprehensive exploration of all possible transformations, the system will impose a punishment ($\text{reward}=-(\hat{n}-n)$), when the number of predicted transformations $n$ is less than the ground truth $\hat{n}$. This mechanism teaches the model to prioritize attempting potentially correct answers over avoiding possible errors, thereby promoting continuous experimentation, in-depth analysis, and precise evaluation of each object’s attribute changes, ultimately leading to steady improvements in accuracy. The punishment reward described above can be formally defined as:
$R^{\text{pun}}=\begin{cases}-n_{\text{mis}}-(\hat{n}-n),&\text{if $n$ < $\hat{n}$};\\ -n_{\text{mis}},&\text{otherwise.}\end{cases}$ (3)
Finally, we sum $R^{\text{pos}}$ and $R^{\text{pun}}$ to obtain $R^{\text{acc}}=R^{\text{pos}}+R^{\text{pun}}$.
3.3 The Learning Algorithm
We employ GRPO as the reinforcement learning algorithm for STAR-R1, since GRPO samples and scores a group of responses based on the reward function, then calculates each response’s relative advantage within the group. Compared to traditional Proximal Policy Optimization (PPO), GRPO eliminates the need for training an additional critic model by estimating the baseline from group scores, thereby significantly improving training efficiency.
Formally, given an initial image $I^{\text{initial}}$, a final image $I^{\text{final}}$, and a question description $Q$, the model first generates a group of $G$ responses $\{T_{1},T_{2},\ldots,T_{G}\}$. STAR-R1 then computes the format reward $R_{g}^{\text{format}}$ and accuracy reward $R_{g}^{\text{acc}}$ for each response, and sums them to obtain a set of rewards $\{R_{1},R_{2},\ldots,R_{G}\}$, where $R_{g}=R_{g}^{\text{format}}+R_{g}^{\text{acc}}$. The advantage $\hat{A}_{g}$ for each response is then calculated by normalizing the rewards with respect to the group statistics: $\hat{A}_{g}=\frac{R_{g}-\mu_{G}}{\sigma_{G}}$, where $\mu_{G}$ and $\sigma_{G}$ are the mean and the standard deviation of the rewards in the group.
Finally, based on the computed advantages and a KL penalty term that constrains the divergence between the trained policy and the reference policy, we obtain the following training objective, which is maximized to optimize the policy model:
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\{T_{g}\}_{g=1}^{G}\sim\pi_{\theta_{old}}(T|q)}$
$\quad\frac{1}{G}\sum_{g=1}^{G}\frac{1}{|o_{i}|}\left[\min\left(R_{g}\hat{A}_{g},\text{clip}\left(R_{g},1-\epsilon,1+\epsilon\right)\hat{A}_{g}\right)-\beta D_{KL}[\pi_{\theta}\|\pi_{ref}]\right],$ (4)
where $r_{g}=\frac{\pi_{\theta}(T_{g}|q)}{\pi_{\theta_{old}}(T_{g}|q)}$ is the probability ratio for observation $T_{g}$ under the current policy $\pi_{\theta}$ relative to the old policy $\pi_{\theta_{old}}$, $\pi_{ref}$ denotes the reference policy, and $\beta$ is a hyperparameter controlling the strength of the KL penalty.
4 Experiments
Dataset. As a challenging multimodal reasoning task, TVR requires the model to infer and describe single-step or multi-step transformations (i.e., changes in object attributes) between given initial and final images. To better reflect real-world scenarios, in addition to simpler tasks where both initial and final images share the same central viewpoint, TVR also includes more challenging cases where the model must match corresponding objects across different viewpoints (e.g., initial image from the central view while final image from the left/right view) and identify the transformations. After cleaning and filtering the original TRANCE dataset (hong2021transformation, ), we randomly sample 9k training examples with both initial and final images from the central view. From the remaining data, we randomly sample 4.5k test examples comprising initial images from the central view, paired with final images from left, center, and right views. We provide further details in Appendix A.
Implementations and Baselines. We employ Qwen2.5-VL-7B as the base model for our main experiments, and conduct all experiments on a single computation node equipped with 8×H20 GPUs. Unlike existing approaches (feng2025video, ; meng2025mm, ) that require multi-stage training, our STAR-R1 demonstrates excellent reasoning capability through direct RL training without any supervised fine-tuning (SFT) as prerequisite. Due to current computational constraints, all RL experiments in this paper (unless otherwise specified) are trained for 2 epochs. To comprehensively evaluate the reasoning capabilities of STAR-R1, we conduct extensive comparisons with numerous existing models, including both closed-source (e.g., Qwen-VL (bai2023qwenvl, ; wang2024qwen2vl, ; bai2025qwen25vl, ), InterVL (chen2024internvl, ; chen2024internvl1.5, ; team2024internvl2, )) and open-source (e.g., GPT-4o (hurst2024gpt, ), Geminiteam2024gemini ) alternatives.
To obtain a comprehensive, intuitive, and in-depth understanding of the model’s reasoning capabilities, we employ multiple evaluation metrics. We first execute all transformations from the model’s response on the initial image state, obtaining the predicted final state $S^{\text{final}}$. By comparing this with the ground truth final state $\hat{S}^{\text{final}}$, we compute the following two categories of metrics:
  • Sample-level Metrics. (1) Attribute Accuracy: measuring whether each attribute (color, shape, size, material) of all objects is predicted correctly. (2) Total Accuracy (TAcc): measuring whether all four attributes of all transformations are predicted correctly. (3) Attribute Difference (Diff): counting the number of differing attributes between $S^{\text{final}}$ and $\hat{S}^{\text{final}}$. (4) Normalized Difference (Norm Diff): dividing Diff by the total number of transformations, eliminating scale effect.
[…]
4.1 Main results
As shown in  Table 1, our model STAR-R1 achieves state-of-the-art performance across all eleven evaluation metrics when compared with advanced MLLMs. This demonstrates both the exceptional capability of STAR-R1 and the remarkable potential of the reinforcement learning paradigm for visual reasoning tasks. Specifically, we have the following observations: (1) For all metrics, STAR-R1 shows significant improvement over the base model Qwen-2.5-VL-7B. It substantially outperforms not only InternVL2 with comparable model scale, but also larger-scale models like QWen-2.5-VL-7B. Remarkably, in terms of TAcc, STAR-R1 surpasses the closed-source commercial models GPT-4o and Gemini-1.5-pro by 37.9% and 45.5%, respectively. (2) Compared to the SFT paradigm, our pure RL approach without any cold-start data achieves approximately 13% performance gain of TAcc, providing strong evidence that RL can significantly enhance model capabilities for complex visual reasoning scenarios. (3) For the four Attribute Accuracy metrics, we observe comparable performance across all attributes. This indicates robust stability in our learning framework without exhibiting catastrophic forgetting or attribute dominance. (4) For the four population-level metrics, model performance shows a steep decline as the number of objects increases. This trend reflects the compounding challenges in complex scenes, where difficulties accumulate across the entire reasoning pipeline from object identification to object association and finally to change perception.
Table 1: Results on TVR task. Bold indicates the best results.
[…]
 GPT-4o hurst2024gpt - 23.5 1.96 0.76 61.7 69.3 62.2 61.0 49.3 27.3 14.6 9.6
Gemini-2.0-flash team2024gemini - 20.9 1.97 0.79 63.3 73.3 59.8 58.0 42.4 24.5 16.8 6.0
Gemini-1.5-Pro team2024gemini - 15.9 2.46 1.00 54.4 72.2 54.6 48.9 37.4 18.5 8.6 4.7
[…]
 Deepseek-VL lu2024deepseek 7B 0.9 5.02 2.23 18.8 38.6 36.4 33.2 4.1 0.1 0.1 0.0
Phi3.5V abdin2024phi 7B 0.9 4.91 2.15 20.1 37.9 39.2 41.1 4.8 0.2 0.1 0.0
mPLUG-owl3 ye2024mplug 7B 1.0 4.82 2.14 19.6 39.1 40.3 40.7 5.1 0.1 0.1 0.0
MiniCPM-V2.6 yao2024minicpm-v 7B 1.0 4.78 2.11 22.5 40.2 41.3 38.3 5.4 0.1 0.1 0.1
LLaVA-OneVision li2024llava-onevision 7B 1.1 4.67 2.06 22.4 43.3 42.5 40.6 5.8 0.1 0.1 0.0
Pixtral agrawal2024pixtral 12B 3.0 3.02 1.44 30.3 53.6 54.7 52.2 10.8 1.9 0.7 0.3
InternVL2.5 chen2024internvl1.5 8B 3.3 2.80 1.14 33.6 59.3 56.0 51.0 11.2 2.7 0.9 0.4
InternVL3 zhu2025internvl3 7B 3.7 2.62 1.08 35.3 58.2 58.5 54.7 13.3 2.6 0.9 0.4
Qwen2-VL wang2024qwen2vl 2B 0.3 5.23 2.40 14.9 45.0 48.4 54.1 1.0 0.3 0.0 0.0
Qwen2-VL wang2024qwen2vl 7B 1.2 4.11 1.84 23.8 45.6 45.3 42.0 6.2 0.1 0.1 0.1
Qwen2.5-VL bai2025qwen25vl 3B 2.3 3.53 1.51 25.6 47.2 52.2 44.5 8.8 1.5 0.5 0.3
Qwen2.5-VL bai2025qwen25vl 7B 3.8 3.41 1.46 26.2 56.5 52.8 45.9 14.5 2.5 0.9 0.3
Qwen2.5-VL bai2025qwen25vl 32B 9.1 3.21 1.38 41.1 57.3 50.6 36.6 27.5 9.8 2.5 1.4
 STAR-SFT 7B 48.7 1.17 0.57 75.5 72.1 80.2 79.3 82.4 53.3 37.9 30.0
STAR-R1 7B 61.4 0.77 0.31 81.3 83.2 86.1 85.5 91.0 70.7 54.2 37.5
[…]
4.2 Anthropomorphic Behavior: RL VS SFT
In this section, we conduct a thorough comparison between SFT and RL training paradigms by evaluating their performance on both In-Domain (ID) and Out-Of-Domain (OOD) test samples, where ID samples contain initial and final images from the same center viewpoint while OOD samples contain images from different viewpoints (center-to-left/right). The OOD setting presents a greater challenge, as models must distinguish genuine object transformations from viewpoint-induced variations. As shown in  Table 2, SFT achieves better accuracy on ID data, but RL demonstrates remarkable superiority in the OOD setting, outperforming SFT by 23%, highlighting its exceptional generalization capability for reasoning in unseen views.
Table 2: Results in In-Domain (ID) and Out-Of-Domain (OOD) scenarios.
[…]
Method TAcc Diff NDiff TAcc Diff NDiff
 STAR-SFT 84.2 0.22 0.08 30.9 1.65 0.83
STAR-R1 76.3 0.38 0.14 53.9 0.96 0.39
[…]
Further case studies reveal that RL model’s superior performance in the OOD scenario stems fundamentally from its capability to systematically compare each object between initial and final images. As illustrated in  Fig. 3, the STAR-SFT model typically performs cursory comparisons of only a few objects. This approach suffices for ID scenarios with consistent viewpoints, where differences are often immediately apparent. However, in OOD scenarios, viewpoint transitions necessitate cross-verification through multiple object comparisons to reliably identify transformations, mirroring human reasoning processes. For example, STAR-SFT misidentifies objects 1, 4, and 3 in the final image as objects 4, 2, and 1, respectively. This suggests that the model fails to effectively account for viewpoint-induced object displacement, erroneously retaining the object IDs from the initial viewpoint in the transformed viewpoint, thus leading to incorrect object correspondence and a significant degradation in prediction accuracy. In contrast, STAR-R1 meticulously examine all objects, including untransformed ones. In multi-object scenes, this comprehensive comparison enables precise cross-view mapping by referring to untransformed objects, thereby preventing performance degradation from mismatches and significantly improving robustness to viewpoint changes. Importantly, during the whole process, our prompts do not include any information related to view invariance/variation. Quantitatively, while processing consistent-viewpoint scenes (ID), RL model performs full object comparisons in 67% of cases, but this ratio increases to 81% for viewpoint-shifting scenarios (OOD). This phenomenon indicates that the model, like humans, adapts its verification intensity based on situational complexity, performing more exhaustive comparisons when viewpoint changes demand higher confidence in object correspondences. Additional case studies are provided in Appendix D for further reference. Furthermore, we apply RL training to the STAR-SFT model and observe a series of intriguing experimental results and phenomena, with the findings presented in Appendix C.
Figure 3: A case study comparing the reasoning processes of STAR-SFT and STAR-R1. STAR-SFT performs only cursory comparisons, failing to detect view changes, which results in incorrect object matching and answers. In contrast, STAR-R1 systematically compares all objects step-by-step, identifies the view shift, and accurately tracks object correspondences, producing correct results. Blue indicates correct answers while red denotes incorrect ones.
4.3 Training Curve of Response Length
We analyze the training curve of the response length to further reveal the underlying insights. For additional details regarding the training curves, please refer to Appendix B. In  Fig. 4, the model’s response length initially undergoes a rapid decline, followed by a gradual increase before eventually stabilizing. We observe that during the early phase, up until the point where the response length reaches its minimum, the model actively explores various reasoning strategies while progressively refining its language. For instance, the model’s output transitions from verbose and multi-object descriptions, such as: "The object with index 3 in the first image is a gray cylinder of size ’large’. In the second image, it has changed to metallic gold. This suggests a transformation in color and material," to
Figure 4: Dynamics of model response length. Zoom in to examine specific examples.
more concise and single-object reasoning, such as: "The small brown sphere changed its color to gray." At this stage, however, the model realizes that focusing on only a few objects may lead to incorrect mappings between objects across the two images or the omission of critical information, leaving room for further improvement in prediction accuracy. Consequently, while maintaining its refined reasoning style for individual objects, the model gradually explores a new and stable reasoning path, systematically comparing the states of all objects between the two images. This approach ultimately achieves high accuracy while keeping the reasoning length at an appropriate level.
4.4 Ablation Studies
In this section, to comprehensively investigate the impact of various factors on model performance, we conduct ablation studies on each component of the reward function, the volume of training data, the maximum response length limit, and the model size.
Reward Design We conducted systematic ablation studies on all components of the reward function in STAR-R1 to evaluate their contributions to model performance, with experimental results presented in Table 4. Our key observations are: (1) The model’s performance degrades when either the object reward (Row 1) or attribute reward (Row 2) is removed, demonstrating that dense rewards facilitate more efficient exploration and improve answer accuracy. (2) Removing the punishment constraint for under-prediction of transformations (when $N_{pred}$ < $N_{gt}$) leads to significant accuracy deterioration (Row 3). This reveals that the punishment mechanism, which encourages exploration and prevents model inaction due to risk aversion, enables the model to search for better reasoning paths. (3) Eliminating the wrong-answer punishment causes the model to reward hacking by exhaustively enumerating all possible (object, attribute, value) triplets, leading to meaningless responses. To further validate the importance of the punishment design, we replace all punishment constraints with a unified negative absolute difference between predicted and ground-truth transformation counts: ($-|n-\hat{n}|$). As shown in Row 4, this yields markedly worse performance, confirming that timely punishment for incorrect answers is crucial for correcting the model’s exploration path. We additionally conduct experiments with Naive GRPO (Row 5), which demonstrates limited suitability for TVR tasks, yielding merely subpar results.
Table 3: Ablation studies on reward design.
[…]
 w/o obj 58.0 0.90 0.37 89.8 66.5 49.5 34.6
w/o attr 56.8 0.96 0.40 87.4 64.6 49.3 34.1
w/o up 58.2 0.97 0.41 90.2 68.5 49.6 33.0
w/o pun 54.3 1.05 0.44 85.5 61.9 46.8 31.4
w naive grpo 54.5 1.02 0.43 85.2 63.7 46.1 31.1
STAR-R1 61.4 0.77 0.31 91.0 70.7 54.2 37.5
[…]
Table 4: Ablation studies on data volume.
[…]
 1,000 13.8 2.32 0.95 29.0 16.2 9.0 4.9
3,000 23.7 1.97 0.88 49.0 25.3 17.2 10.3
5,000 40.7 1.49 0.65 75.1 47.0 30.1 19.5
7,000 52.6 1.08 0.46 82.7 59.6 45.0 31.2
9,000 61.4 0.77 0.31 91.0 70.7 54.2 37.5
[…]
Training Data To evaluate the effect of training sizes, we randomly sample four subsets for training from the whole training set, containing 1000, 3000, 5000, and 7000 samples, respectively, and evaluate them on the original test set. The experimental results shown in  Table 4 demonstrate that as the volume of training data increases, model performance improves continuously, but the growth rate of accuracy exhibits an initial increase followed by a subsequent decline. This suggests that insufficient data in the early stages constrains model learning, while expanded datasets facilitate a phase of rapid knowledge acquisition, leading to significantly accelerated performance gains.
Table 5: Ablation studies on model size.
[…]
 Qwen2.5-VL-3B 2.3 3.53 1.51 8.8 1.5 0.5 0.3
STAR-R1-3B 28.6 1.84 0.79 55.2 32.4 20.4 13.5
 Qwen2.5-VL-7B 3.8 3.41 1.46 14.5 2.5 0.9 0.3
STAR-R1-7B 61.4 0.77 0.31 91.0 70.7 54.2 37.5
[…]
Model Size To assess the influence of model size, we perform experiments by replacing the base model with Qwen-2.5VL-3B. The experimental results are presented in  Table 5. It is evident that the 3B model also exhibits a significant improvement in prediction accuracy after RL training, demonstrating the effectiveness of our method. However, under the same training settings, the performance gain of the 3B model is considerably smaller than that of the 7B model. This suggests that the capability of the base model plays a crucial role in the post-training phase, determining the model’s ability to learn and adapt to new knowledge and scenarios. A stronger base model can achieve a higher upper bound of reasoning performance.
5 Conclusion
In this paper, we present STAR-R1, a pure RL-trained model equipped with our new reward design for the challenging spatial reasoning task—TVR. By integrating dense rewards for partial correctness with punishment rewards for penalizing shortcuts, STAR-R1 enables robust learning, achieving superior performance across all eleven evaluation metrics. Results reveal a mechanistic advantage of RL over SFT stems from comprehensively comparing objects in the scene, especially in cross-view out-of-domain scenarios. Further analysis uncovers STAR-R1’s anthropomorphic behavior and non-linear training dynamic of response length, offering insights into RL’s capability in complex visual reasoning. This work validates the potential of the R1-Zero paradigm for advanced reasoning tasks. Our findings may pave the way for future research in enhancing MLLMs’ spatial cognition, emphasizing RL’s capability to unlock human-like reasoning through structured exploration and fine-grained feedback.